# Dual Magnetic and Electric Dipole Symmetry: Pseudo-Angular Momentum in Parity Space and the Electric Landé $g$-Factor

### v1

@@ V The Electric G-factor
Intrinsic operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator $\hat{\vec{S}}=\frac{\hbar}{2}\bm{\sigma}$ Spin operator
@@ VI Discussion and Implications
The Stark–sector theory developed here is the quantum dual of that classical picture. Starting from the wavefunction’s microscopic polarization $\vec{P}(\vec{r})$, a nonzero curl defines an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$, in a similar way to the classical picture [12, 13, 14, 25]. Thus, the induced EDM in a static field is governed by a circulating magnetic probability current in the orbital (Runge-Lenz) sector, mirroring how an impressed electric field in classical circuit and antenna theory is traded for an equivalent magnetic surface current that generates the same exterior fields. In this sense the dual-Ohanian formulation makes explicit the classical-quantum correspondence, whereby an impressed electric drive is represented by a magnetic equivalence current in the classical dipole [12, 13], while Stark mixing is represented by a magnetic probability current in the quantum dipole.

### v2

@@ V The Electric G-factor
Intrinsic operator $\hat{\vec{S}}=\frac{\hbar}{2}\vec{\sigma}$ Spin operator $\hat{\vec{S}}=\frac{\hbar}{2}\vec{\sigma}$ Spin operator
@@ VI Discussion and Implications
The Stark-sector theory developed here is the quantum dual of the classical equivalence-current picture. Starting from the wavefunction’s microscopic polarization $\vec{P}(\vec{r})$, we define an effective magnetic probability current, $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$, which provides a dual-current representation of the polarization texture in close analogy with the classical construction [12, 13, 14, 27].
Importantly, the induced EDM in a static field arises physically from charge displacement via parity (Stark) mixing; $\vec{J}_{m}$ is not an additional causal mechanism, but a convenient field-equivalent description that encodes the same orbital (Runge-Lenz) structure. This mirrors how, in classical circuit and antenna theory, an impressed electric drive can be traded for an equivalent magnetic surface current that reproduces the same exterior fields. In this sense, the dual-Ohanian formulation makes the classical-quantum correspondence explicit: an impressed electric excitation is represented by a magnetic equivalence current in the classical dipole [12, 13], while Stark mixing is represented by the magnetic probability-current representation of the quantum dipole.
@@ Appendix A Runge-Lenz Operator
Appendix C Hydrogenic $SO(4)$ structure and the pseudo-angular momentum
For a hydrogenic atom/ion, the following familiar Coulomb-problem commutation relations encode the hidden $SO(4)$ symmetry of the bound-state ($H<0$) spectrum.
$[L_{i},L_{j}]$ $=i\hbar\,\epsilon_{ijk}L_{k},$ (78)
$[L_{i},A_{j}]$ $=i\hbar\,\epsilon_{ijk}A_{k},$ (79)
$[A_{i},A_{j}]$ $=-2\mu H\,i\hbar\,\epsilon_{ijk}L_{k},$ (80)
For bound states it is convenient to use the scaled Runge–Lenz operator (75) , so (80) implies
$A_{{\rm sc},i},A_{{\rm sc},j}]\;=\;i\hbar\,\epsilon_{ijk}L_{k},$ (81)
and $\vec{A}_{\rm sc}$ has the dimensions of angular momentum (like $\vec{L}$). One may then form the standard $SO(4)\simeq SU(2)\times SU(2)$ generators
$\vec{J}_{\pm}\;\equiv\;\frac{1}{2}\left(\vec{L}\pm\vec{A}_{\rm sc}\right),$ (82)
which satisfy the $SU(2)$ algebras
$[J_{\pm,i},J_{\pm,j}]$ $=i\hbar\,\epsilon_{ijk}J_{\pm,k},$ (83)
$[J_{+,i},J_{-,j}]$ $=0.$ (84)
Within a fixed $n$-manifold (degenerate subspace), these relations provide the rigorous $SU(2)$ structure underpinning the Landé-factor analogy used below.
In the Stark/parity (two- or few-level) subspace relevant for the orbital EDM, we identify the pseudo-angular momentum $\vec{J}_{p}$ with the appropriate projected $SU(2)$ generator (equivalently, the corresponding component of $\vec{J}_{\pm}$ along the field axis) that labels the parabolic/Stark structure. Consequently, $\vec{J}_{p}$ obeys the standard $SU(2)$ commutation relations,
$J_{p,i},J_{p,j}]\;=\;i\hbar\,\epsilon_{ijk}J_{p,k},$ (85)
justifying the Landé-factor form $\hat{\vec{d}}_{\rm orb}=g_{E}d_{B}\,\hat{\vec{J}}_{p}/\hbar$ used in this work.